OpenStudy (anonymous):

Given the following endpoint (A) and midpoint (M), find the other endpoint (B). M(2, 3) A(0, -1) 1) (-1, 0) 2) (7, 4) 3) (4, 7) 4) (3, 2)

OpenStudy (amistre64):

(a+b)/2 = m 0,-1 +x , y ------ (0+x, -1+y) (0+x, -1+y)/2 = ( 2,3) (0+x)/2 = 2 ; (-1+y)/2 = 3 0+x = 4 -1+y = 6 x = 4 y = 7 ( 4 , 7 ) id say option 3